# 00footnotetext: Shared first author; Shared last author

### v1

@@ Abstract
We introduce DataComp for Language Models (DCLM), a testbed for controlled dataset experiments with the goal of improving language models. As part of DCLM, we provide a standardized corpus of 240T tokens extracted from Common Crawl, effective pretraining recipes based on the OpenLM framework, and a broad suite of 53 downstream evaluations. Participants in the DCLM benchmark can experiment with data curation strategies such as deduplication, filtering, and data mixing at model scales ranging from 412M to 7B parameters. As a baseline for DCLM, we conduct extensive experiments and find that model-based filtering is key to assembling a high-quality training set. The resulting dataset, DCLM-baseline, enables training a 7B parameter language model from scratch to 6464% 5-shot accuracy on MMLU with 2.6T training tokens. Compared to MAP-Neo, the previous state-of-the-art in open-data language models, DCLM-baseline represents a 6.6 percentage point improvement on MMLU while being trained with 40% less compute. Our baseline model is also comparable to Mistral-7B-v0.3 and Llama 3 8B on MMLU (63% & 66%), and performs similarly on an average of 53 natural language understanding tasks while being trained with $6.6\times$ less compute than Llama 3 8B. Our results highlight the importance of dataset design for training language models and offer a starting point for further research on data curation.
@@ Introduction
Figure 1: Improving training sets leads to better models that are cheaper to train. Using DataComp-LM, we develop a high-quality dataset, DCLM-baseline, which we use to train models with state-of-the-art trade-off between compute and performance. We compare on both (left) a Core set of tasks and on (right) MMLU 5-shot. Specifically DCLM-baseline (orange) shows favorable performance relative to both close-source models (crosses) and other open-source datasets and models (circles).
@@
The high cost of training language models makes it necessary to understand the performance of training recipes across different compute and data scales. Hence, our third contribution is an investigation of scaling trends for dataset design. We find that models as small as 400M parameters can still provide signal on which training sets perform better at larger scales. Based on our experiments, we organize DCLM into five compute scales spanning a range of about $600\times$ in compute from 400M parameter models to over-trained 7B models. This multi-scale design makes DCLM accessible to researchers with varying compute budgets.
@@
Finally, we combine our results into DCLM-baseline, a new state-of-the-art public training set for language models. When training a 7B parameter language model on 2.6 trillion tokens using DCLM-baseline, the resulting model achieves 6464% on MMLU, which is state-of-the-art among open-data models and close to models such as Mistral-7B-v0.3 (63%) or Llama 3 8B (66%) that are trained with up to $6.6\times$ more compute (Llama 3 8B). Compared to Llama 2 7B, training a 7B parameter model on 280B tokens from DCLM-baseline achieves 5 pp higher MMLU while being trained with $7\times$ less compute. As our 7B model uses a standard decoder-only Transformer [161, 124, 157], our results also highlight that a systematic approach to data curation is key to training performant language models.
We publicly release our DCLM infrastructure, models, and training sets at https://datacomp.ai/dclm to enable other researchers to participate in DCLM and to strengthen the empirical foundations for data-centric research on language models.
@@
We summarize closely related work in this section and provide additional related work in Appendix˜B.
@@
As the scale of LMs has increased over the past years [157, 158, 4, 34, 71, 112, 125, 151], the community has curated larger datasets to match. Early works include the C4 dataset with 160 billion (B) tokens and The Pile [57] with 300B tokens. More recently, RefinedWeb [118] contains 600B tokens, Dolma [147] 3 trillion (T) tokens, FineWeb 15T tokens [119], and RedPajama-v2 30T tokens [41]. There are also large domain-specific datasets, such as the code-focused StackV2 with 900B tokens [98], as well as high-quality filtered subsets such as FineWeb-Edu [97] with 1.3T tokens. We include performance comparisons with various datasets in Figure˜1 and examine FineWeb’s LightEval evaluation framework more closely in Appendix˜G. We release the largest pool of raw text data to date with 240T web-crawled tokens. We also release DCLM-baseline, a high-quality dataset from our pool that yields better models than prior datasets.
@@
Table 1: DCLM competition scales. DCLM contains five competition scales, enabling research in varying compute regimes. Each scale specifies the model size (‘Model parameters’, NN), the number of tokens seen during training (‘Train tokens’, $D$), and the size of the original pool that can be used for filtering (‘Pool size’). We provide an estimate of the compute required for training (‘Train FLOPs’=6ND=6ND) and GPU hours (‘Train H100 hours’) using the OpenLM training framework [68].
@@
This section describes the main components of DCLM. We start with DCLM-Pool, the raw text corpus underlying our benchmark (Section˜3.1). We then develop the DCLM workflow, visualized in Figure˜2: selecting a competition scale (Section˜3.2), curating a dataset by filtering DCLM-Pool and potentially mixing in other sources (Section˜3.3), training a model with fixed hyperparameters (Section˜3.4), and evaluating the model to score the dataset (Section˜3.5).
@@
DCLM-Pool is an unfiltered web-text corpus comprised of all Common Crawl [40] data prior to 2023. Based on Section˜4.2, we re-extract text from HTML using resiliparse [19, 20] instead of using Common Crawl’s pre-extracted text. DCLM-Pool contains 200B documents (370TB after gzip compression), resulting in 240T GPT-NeoX [22] tokens. See Appendix˜E for additional details.
@@
Test set samples often contaminate language model training sets [46, 177, 49]; however, the effect of such samples on downstream performance remains largely unclear [86, 147, 112]. To allow researchers to better understand contamination, we release decontamination tooling instead of decontaminating DCLM-Pool directly. Our tools are based on Lee et al. [86] and allow participants to examine their datasets for overlap with our test sets. We ask all submissions to disclose a decontamination report and avoid using highly-contaminated data. For the highest scoring submissions, we plan to specifically evaluate them for contamination. In Section˜4.6, we apply our tools to DCLM-Pool and evaluate whether contamination affects our models.
@@
To ensure DCLM is accessible to researchers with different compute constraints and to facilitate the study of scaling trends, we create different competition scales spanning three orders of compute magnitude (Table˜1).Each scale (i.e., 400M-1x, 1B-1x, 1B-5x, 7B-1x, and 7B-2x) specifies the number of model parameters (e.g., 7B) and a Chinchilla multiplier (e.g., 1x). The number of training tokens for each scale is 2020 $\times$ number of parameters $\times$ Chinchilla multiplier so that a multiplier of 1x corresponds to a compute allocation that Hoffmann et al. [71] found near-optimal.
A potential pitfall in our multi-scale design is that the ranking of data curation methods may change when increasing the compute scale. To better understand this concern, in Figure˜3, we plot the performance of 10 methods at the 7B-1x scale as a function of their 400M-1x and 1B-1x performance. We find high rank correlation between the smaller 400M-1x, 1B-1x results and larger 7B-1x results (Pearson’s $r=0.885$ and $r=0.919$, respectively), suggesting better curation strategies at smaller scales transfer to larger scales. For more competition scale ablations, including experiments suggesting dataset improvements are largely orthogonal to training hyperparameters, see Appendix˜H.
@@
After choosing a scale, participants choose one of two tracks. (i) In the filtering track, participants propose algorithms to select training data from a candidate pool. We start with five pools, one for each scale in Table˜1, which are random document subsets of DCLM-Pool. We restrict initial pool sizes by scale to encourage scalable filtering strategies and reflect realistic data download and storage constraints. (ii) In the mixing track, a submission combines documents from potentially many sources. For instance, participants can synthesize documents from DCLM-Pool, a custom crawl, Stack Overflow, and Wikipedia. Appendix˜C provides detailed rules for each track, and Appendix˜D describes our open-source, extensible tooling for executing filtering and mixing operations.
@@
To isolate the effect of dataset interventions, we fix a training recipe at each scale. Based on prior ablations on model architectures and training [124, 170, 56, 85, 71, 28, 34, 157, 158, 4], we adopt a decoder-only Transformer (e.g., GPT-2, Llama) [161, 157, 124], implemented in OpenLM [68]. We also provide unified data processing utilities. Appendix˜F contains additional training details.
@@
Our full evaluation suite, based on LLM-Foundry [106], contains 53 downstream tasks suitable for base model evaluation (i.e., without fine-tuning): from question answering to open-ended generation formats, considering varied domains like coding, text-book knowledge, and common-sense reasoning. To evaluate data curation algorithms, we focus on three main performance metrics. First, we consider MMLU 5-shot accuracy [70], which is widely used to compare state-of-the-art models like GPT-4 [112] and Llama 3 70B [4]. Second, we propose Core centered accuracy, computed over a subset of 22 tasks (e.g., HellaSwag [182] and ARC-E [38]) that provide a low-variance signal even at small scales, linearly rescaling the accuracy per task so that 0 corresponds to random guessing and 1 corresponds to perfect accuracy. Finally, we report Extended centered accuracy, which averages the centered performance for all of our 53 tasks. For more metric details, see Appendix˜G.
@@
We now show how the DCLM workflow can lead to high-quality datasets and quantify the effect of data curation methods. This section describes the process of converting Common Crawl into our dataset, DCLM-baseline, as shown in Figure˜4. We provide ablation experiments for each step along the way. We first evaluate open-source datasets as a starting point (Section˜4.1). Next, we experiment with alternatives for several key phases of dataset construction: text extraction (Section˜4.2), deduplication (Section˜4.3), and model-based filtering (Section˜4.4). We then experiment with mixing in high-quality sources (Section˜4.5) and provide a contamination analysis (Section˜4.6). In Section˜5, we scale up this approach to train a 7B model for 2T tokens.
@@
We begin by evaluating several well-known open-source datasets (C4 [127, 46], RefinedWeb [118], RedPajama [156], and Dolma-V1 [147]) in Table˜2. While all four datasets use various heuristic filters and data cleaning steps, we find that RefinedWeb performs the best on our Core and Extended metrics at the 7B-1x scale. RefinedWeb applies the following filtering pipeline: Common Crawl text extraction, heuristic selection rules (e.g., to remove spam), and deduplication of repeated content. Interestingly, RefinedWeb is solely filtered from Common Crawl, unlike RedPajama and Dolma-V1, which additionally mix in curated, “high-quality” sources like Wikipedia. The comparison suggests the relative strength of filtering, which we explore later in our experiments.
@@
Text extraction is a common early processing step that pulls content from raw HTML. To understand the effect of this step, we compare three text extraction approaches: resiliparse, trafilatura (used by RefinedWeb), and the Common Crawl-provided WET files that contain pre-extracted text. We then apply RefinedWeb’s heuristic quality filters to each of the text extractions. In Table˜3, we find both resiliparse and trafilatura improve Core by at least 2.5 points over the WET extraction. This is significant because most open source datasets, including C4, RedPajama, and Dolma-V1, use the WET extraction, which could partially explain their worse performance in Table˜2. While resiliparse and trafilatura have similar downstream performance, resiliparse is $8\times$ faster to run and hence more practical for large-scale processing. For more analysis, see Appendix˜J.
@@
Web-crawled datasets often contain many duplicate or near-duplicate data strings. Removing these duplicates from a training set serves the dual purpose of improving performance by reducing memorization [31, 86] and increasing data diversity. For deduplication, we explore MinHash [26], as part of a suffix array pipeline [86, 118], and near-duplicate Bloom filtering, which modifies an exact document and paragraph deduplication scheme [147]. We find that both approaches provide comparable downstream performance: within 0.2 Core percentage points at the 7B-2x scale. However, our modified Bloom filter approach scales more easily to datasets surpassing 10TB. We provide additional analysis in Appendix˜K.
@@
We compare many strategies: 1) PageRank score filtering to retain documents based on how likely they are to be linked to other documents, 2) Semantic Deduplication (SemDedup) to remove documents with similar informational content [1], 3) linear classifiers fit on pre-trained BGE text embeddings [172], 4) AskLLM that prompts an LM to see if a document is helpful [136], 5) Perplexity filtering where we retain low perplexity sequences following CCNet [166], 6) Top-k average logits where we average the top-$k$ model logits over all words in a document to score how confident a model is that the correct words are within $k$ reasonable choices, and 7) fastText [79] binary classifiers to distinguish data quality. For training classifiers, we train on $\sim 400$k documents split equally between positive and negative classes. We experiment with different options for positive data and fix negative data as a sample from RefinedWeb. For the perplexity filtering and the top-k average logits strategies, we utilize a 154M parameter causal Transformer trained on a mix of English Wikipedia, the books subset of RedPajama v1, and peS2o [146, 156]. We compare the aforementioned approaches in Table˜4 and find that fastText-based filtering outperforms all other approaches. We next aim to understand how fastText training recipes affect its effectiveness as a data filtering network [53].
@@
To better understand the limits of fastText, we train several variants, exploring different choices for the reference data (i.e., the examples given positive labels), feature space, and filtering threshold, as shown in Table˜5. For reference positive data, we considered commonly used sources like Wikipedia [57], OpenWebText2 [57], and RedPajama-books [156], following the reference data used for GPT-3 [28]. We also try a novel approach, using instruction-formatted data, drawing examples from OpenHermes 2.5 [153] (OH-2.5) and high-scoring posts from the r/ExplainLikeImFive (ELI5) subreddit. Overall, we find, when controlling for other hyperparameters, the fastText OH-2.5 +ELI5 approach gives a 3.5 percentage point lift on Core compared to the conventional choices. It is natural to ask whether using OH-2.5 data for filtering could preclude additional gains from instruction-tuning. In Appendix˜P, we show this is not the case, further suggesting the strength and compatibility of this approach with modern fine-tuning paradigms. Finally, we observe that using a fairly strict threshold, which keeps the top-10% of examples, helps over more permissive top-15% and top-20% thresholds. We further study the unintuitive behavior of dataset filtering and its connection to human judgment in Appendix˜M.
@@
Researchers often combine Common Crawl (CC) with other data sources that are considered high-quality [156, 57, 63, 158] (e.g., Wikipedia, arXiv, Stack exchange, and peS2o [146]). Since DCLM participants can include additional data sources in the bring your own data track, we examined the potential benefits of adding high-quality sources to training sets derived from Common Crawl only. We compare a model trained on 100% filtered CC data to models trained with the mixing proportion from Llama 1 and RedPajama: 67% CC, and 33% from Wikipedia, Books, Stack exchange, arXiv, and Github. For the CC component, we consider different variants: a subset of our DCLM-baseline, RedPajama’s CC portion, RefinedWeb, and C4. The results in Table˜6 show that mixing improves performance for the lower-performing CC subsets (C4, RedPajama-CC, and RefinedWeb). In the case of DCLM-baseline however, mixing actually hurts performance on average, which suggests it can be counterproductive given performant filtering. For additional mixing results, see Appendix˜L.
@@
C4 23.7 25.9 (+2.2+2.2) 12.5 13.3 (+0.8+0.8)
RPJ CC only 24.0 25.7 (+1.7+1.7) 12.1 13.5 (+1.4+1.4)
RefinedWeb 25.1 26.5 (+1.4+1.4) 12.9 13.1 (+0.2+0.2)
DCLM-baseline  [^1] The DCLM-baseline subset used here relied on an earlier version of our classifier and our pool. 31.1 29.5 (1.6-1.6) 16.0 15.3 (0.7-0.7)
@@
The results of this analysis can be seen in Table˜7. We see that this removal of contaminated samples does not lead to a decrease in performance for our model. As such, we can see that our performance gains in MMLU are not caused by increased presence of MMLU in our dataset.
@@
We also apply the above removal strategy on Dolma-V1.7 [147] and FineWeb-Edu [97], to measure what the contamination differences are between DCLM-baseline and those datasets. The results can be seen in Table˜8. We see that our DCLM-baseline has roughly similar contamination stats as other high performing datasets according to this analysis. We provide further contamination analysis that extends to our entire benchmark suite in Appendix˜N.
@@
Here, we test if datasets that perform well on the DCLM benchmark also maintain their strength with an order of magnitude more compute. To ensure our trained model is broadly useful, including for math and coding tasks, we combine our 3.8T DCLM-baseline with the StarCoder [88] and ProofPile2 [13] data to arrive at a 4.1T token dataset. We train a 7B model for 2.5T tokens on this dataset with the same hyperparameters as our largest competition scale except for an annealing phase for the 200B tokens on a modified distributed that was 70% DCLM-baseline with a tighter fastText threshold, and 30% math datasets (see Appendix˜P). We then adopt the continual pre-training methodology from Pouransari et al. [123] for 100B tokens on the same distribution to increase the context length from 2048 to 8192, we provide more details on this procedure in Section˜P.2.
In Table˜9, we show that our model outperforms all 7B models trained on public training sets and approaches closed-data models trained for more tokens such as Llama-8B, Mistral-7B, and Gemma-7B. Additionally, in Appendix˜O, we show that our model achieves strong instruction-tuning performance. After instruction tuning on publicly available IT datasets, our model maintains most of its benchmark performance and achieves an AlpacaEval2.0 LC Win-rate of 16.6, which outperforms Gemma-Instruct (10.4), while approaching the strong performance of Mistral-v0.2-7B (17.1) and Llama3-Instruct (22.9).

### v2

@@ Abstract
We introduce DataComp for Language Models (DCLM), a testbed for controlled dataset experiments with the goal of improving language models. As part of DCLM, we provide a standardized corpus of 240T tokens extracted from Common Crawl, effective pretraining recipes based on the OpenLM framework, and a broad suite of 53 downstream evaluations. Participants in the DCLM benchmark can experiment with data curation strategies such as deduplication, filtering, and data mixing at model scales ranging from 412M to 7B parameters. As a baseline for DCLM, we conduct extensive experiments and find that model-based filtering is key to assembling a high-quality training set. The resulting dataset, DCLM-baseline, enables training a 7B parameter language model from scratch to $64$% 5-shot accuracy on MMLU with 2.6T training tokens. Compared to MAP-Neo, the previous state-of-the-art in open-data language models, DCLM-baseline represents a 6.6 percentage point improvement on MMLU while being trained with 40% less compute. Our baseline model is also comparable to Mistral-7B-v0.3 and Llama 3 8B on MMLU (63% & 66%), and performs similarly on an average of 53 natural language understanding tasks while being trained with $\times6$ less compute than Llama 3 8B. Our results highlight the importance of dataset design for training language models and offer a starting point for further research on data curation. We release the DCLM benchmark, framework, models, and datasets at https://datacomp.ai/dclm.
@@ Introduction
Figure 1: Improving training sets leads to better models that are cheaper to train. Using DataComp-LM, we develop a high-quality dataset, DCLM-baseline, which we use to train models with state-of-the-art trade-off between compute and performance. We compare on both (left) a Core set of tasks and on (right) MMLU 5-shot. Specifically DCLM-baseline (orange) shows favorable performance relative to both close-source models (crosses) and other open-source datasets and models (circles). Models in this figure are from [161, 162, 4, 97, 22, 189, 154, 121, 43, 160, 68, 150, 130, 100, 10, 156].
@@
The high cost of training language models makes it necessary to understand the performance of training recipes across different compute and data scales. Hence, our third contribution is an investigation of scaling trends for dataset design. We find that models as small as 400M parameters can still provide signal on which training sets perform better at larger scales. Based on our experiments, we organize DCLM into five compute scales spanning a range of about 600×600$\times$600 × in compute from 400M parameter models to over-trained 7B models. This multi-scale design makes DCLM accessible to researchers with varying compute budgets.
@@
Finally, we combine our results into DCLM-baseline, a new state-of-the-art public training set for language models. When training a 7B parameter language model on 2.6 trillion tokens using DCLM-baseline, the resulting model achieves $64$% on MMLU, which is state-of-the-art among open-data models and close to models such as Mistral-7B-v0.3 (63%) or Llama 3 8B (66%) that are trained with up to $\times6$ more compute (Llama 3 8B). Compared to Llama 2 7B, training a 7B parameter model on 280B tokens from DCLM-baseline achieves 5 pp higher MMLU while being trained with 7×7$\times$7 × less compute. As our 7B model uses a standard decoder-only Transformer [165, 127, 161], our results also highlight that a systematic approach to data curation is key to training performant language models.
We publicly release our DCLM framework, models, and training sets at https://datacomp.ai/dclm to enable other researchers to participate in DCLM and to strengthen the empirical foundations for data-centric research on language models.
@@
We summarize closely related work in this section and provide additional related work in Appendix B.
@@
As the scale of LMs has increased over the past years [161, 162, 4, 36, 73, 115, 128, 154], the community has curated larger datasets to match. Early works include the C4 dataset with 160 billion (B) tokens and The Pile [59] with 300B tokens. More recently, RefinedWeb [121] contains 600B tokens, Dolma [150] 3 trillion (T) tokens, FineWeb 15T tokens [122], and RedPajama-v2 30T tokens [43]. There are also large domain-specific datasets, such as the code-focused StackV2 with 900B tokens [101], as well as high-quality filtered subsets such as FineWeb-Edu [100] with 1.3T tokens. We include performance comparisons with various datasets in Figure 1 and examine FineWeb’s LightEval evaluation framework more closely in Appendix G. We release the largest pool of raw text data to date with 240T web-crawled tokens. We also release DCLM-baseline, a high-quality dataset from our pool that yields better models than prior datasets.
@@
Table 1: DCLM competition scales. DCLM contains five competition scales, enabling research in varying compute regimes. Each scale specifies the model size (‘Model parameters’, $N$), the number of tokens seen during training (‘Train tokens’, $D$), and the size of the original pool that can be used for filtering (‘Pool size’). We provide an estimate of the compute required for training (‘Train FLOPs’$=6ND$) and GPU hours (‘Train H100 hours’) using the OpenLM training framework [70].
@@
This section describes the main components of DCLM. We start with DCLM-Pool, the raw text corpus underlying our benchmark (Section 3.1). We then develop the DCLM workflow, visualized in Figure 2: selecting a competition scale (Section 3.2), curating a dataset by filtering DCLM-Pool and potentially mixing in other sources (Section 3.3), training a model with fixed hyperparameters (Section 3.4), and evaluating the model to score the dataset (Section 3.5).
@@
DCLM-Pool is an unfiltered web-text corpus comprised of all Common Crawl [42] data prior to 2023. Based on Section 4.2, we re-extract text from HTML using resiliparse [20, 21] instead of using Common Crawl’s pre-extracted text. DCLM-Pool contains 200B documents (370TB after gzip compression), resulting in 240T GPT-NeoX [24] tokens. See Appendix E for additional details.
@@
Test set samples often contaminate language model training sets [48, 181, 51]; however, the effect of such samples on downstream performance remains largely unclear [88, 150, 115]. To allow researchers to better understand contamination, we release decontamination tooling instead of decontaminating DCLM-Pool directly. Our tools are based on Lee et al. [88] and allow participants to examine their datasets for overlap with our test sets. We ask all submissions to disclose a decontamination report and avoid using highly-contaminated data. For the highest scoring submissions, we plan to specifically evaluate them for contamination. In Section 4.6, we apply our tools to DCLM-Pool and evaluate whether contamination affects our models.
@@
To ensure DCLM is accessible to researchers with different compute constraints and to facilitate the study of scaling trends, we create different competition scales spanning three orders of compute magnitude (Table 1).Each scale (i.e., 400M-1x, 1B-1x, 1B-5x, 7B-1x, and 7B-2x) specifies the number of model parameters (e.g., 7B) and a Chinchilla multiplier (e.g., 1x). The number of training tokens for each scale is $20$$\times$ number of parameters $\times$ Chinchilla multiplier so that a multiplier of 1x corresponds to a compute allocation that Hoffmann et al. [73] found near-optimal.
A potential pitfall in our multi-scale design is that the ranking of data curation methods may change when increasing the compute scale. To better understand this concern, in Figure 3, we plot the performance of 10 methods at the 7B-1x scale as a function of their 400M-1x and 1B-1x performance. We find high rank correlation between the smaller 400M-1x, 1B-1x results and larger 7B-1x results (Pearson’s $r=0.885$ and $r=0.919$, respectively), suggesting better curation strategies at smaller scales transfer to larger scales. For more competition scale ablations, including experiments suggesting dataset improvements are largely orthogonal to training hyperparameters, see Appendix H.
@@
After choosing a scale, participants choose one of two tracks. (i) In the filtering track, participants propose algorithms to select training data from a candidate pool. We start with five pools, one for each scale in Table 1, which are random document subsets of DCLM-Pool. We restrict initial pool sizes by scale to encourage scalable filtering strategies and reflect realistic data download and storage constraints. (ii) In the mixing track, a submission combines documents from potentially many sources. For instance, participants can synthesize documents from DCLM-Pool, a custom crawl, Stack Overflow, and Wikipedia. Appendix C provides detailed rules for each track, and Appendix D describes our open-source, extensible tooling for executing filtering and mixing operations.
@@
To isolate the effect of dataset interventions, we fix a training recipe at each scale. Based on prior ablations on model architectures and training [127, 174, 58, 87, 73, 30, 36, 161, 162, 4], we adopt a decoder-only Transformer (e.g., GPT-2, Llama) [165, 161, 127], implemented in OpenLM [70]. We also provide unified data processing utilities. Appendix F contains additional training details.
@@
Our full evaluation suite, based on LLM-Foundry [109], contains 53 downstream tasks suitable for base model evaluation (i.e., without fine-tuning): from question answering to open-ended generation formats, considering varied domains like coding, text-book knowledge, and common-sense reasoning. To evaluate data curation algorithms, we focus on three main performance metrics. First, we consider MMLU 5-shot accuracy [72], which is widely used to compare state-of-the-art models like GPT-4 [115] and Llama 3 70B [4]. Second, we propose Core centered accuracy, computed over a subset of 22 tasks (e.g., HellaSwag [186] and ARC-E [40]) that provide a low-variance signal even at small scales, linearly rescaling the accuracy per task so that 0 corresponds to random guessing and 1 corresponds to perfect accuracy. Finally, we report Extended centered accuracy, which averages the centered performance for all of our 53 tasks. For more metric details, see Appendix G.
@@
We now show how the DCLM workflow can lead to high-quality datasets and quantify the effect of data curation methods. This section describes the process of converting Common Crawl into our dataset, DCLM-baseline, as shown in Figure 4. We provide ablation experiments for each step along the way. We first evaluate open-source datasets as a starting point (Section 4.1). Next, we experiment with alternatives for several key phases of dataset construction: text extraction (Section 4.2), deduplication (Section 4.3), and model-based filtering (Section 4.4). We then experiment with mixing in high-quality sources (Section 4.5) and provide a contamination analysis (Section 4.6). In Section 5, we scale up this approach to train a 7B model for 2T tokens.
@@
Evaluating existing training datasets
We begin by evaluating several well-known open-source datasets (C4 [130, 48], RefinedWeb [121], RedPajama [160], and Dolma-V1 [150]) in Table 2. While all four datasets use various heuristic filters and data cleaning steps, we find that RefinedWeb performs the best on our Core and Extended metrics at the 7B-1x scale. RefinedWeb applies the following filtering pipeline: Common Crawl text extraction, heuristic selection rules (e.g., to remove spam), and deduplication of repeated content. Interestingly, RefinedWeb is solely filtered from Common Crawl, unlike RedPajama and Dolma-V1, which additionally mix in curated, “high-quality” sources like Wikipedia. The comparison suggests the relative strength of filtering, which we explore later in our experiments.
@@
Text extraction is a common early processing step that pulls content from raw HTML. To understand the effect of this step, we compare three text extraction approaches: resiliparse, trafilatura (used by RefinedWeb), and the Common Crawl-provided WET files that contain pre-extracted text. We then apply RefinedWeb’s heuristic quality filters to each of the text extractions. In Table 3, we find both resiliparse and trafilatura improve Core by at least 2.5 points over the WET extraction. This is significant because most open source datasets, including C4, RedPajama, and Dolma-V1, use the WET extraction, which could partially explain their worse performance in Table 2. While resiliparse and trafilatura have similar downstream performance, resiliparse is 8×8$\times$8 × faster to run and hence more practical for large-scale processing. For more analysis, see Appendix J.
@@
Web-crawled datasets often contain many duplicate or near-duplicate data strings. Removing these duplicates from a training set serves the dual purpose of improving performance by reducing memorization [33, 88] and increasing data diversity. For deduplication, we explore MinHash [28], as part of a suffix array pipeline [88, 121], and near-duplicate Bloom filtering, which modifies an exact document and paragraph deduplication scheme [150]. We find that both approaches provide comparable downstream performance: within 0.2 Core percentage points at the 7B-2x scale. However, our modified Bloom filter approach scales more easily to datasets surpassing 10TB. We provide additional analysis in Appendix K.
@@
We compare many strategies: 1) PageRank score filtering to retain documents based on how likely they are to be linked to other documents, 2) Semantic Deduplication (SemDedup) to remove documents with similar informational content [1], 3) linear classifiers fit on pre-trained BGE text embeddings [176], 4) AskLLM that prompts an LM to see if a document is helpful [139], 5) Perplexity filtering where we retain low perplexity sequences following CCNet [170], 6) Top-k average logits where we average the top-$k$ model logits over all words in a document to score how confident a model is that the correct words are within $k$ reasonable choices, and 7) fastText [81] binary classifiers to distinguish data quality. For training classifiers, we train on $\sim 400$k documents split equally between positive and negative classes. We experiment with different options for positive data and fix negative data as a sample from RefinedWeb. For the perplexity filtering and the top-k average logits strategies, we utilize a 154M parameter causal Transformer trained on a mix of English Wikipedia, the books subset of RedPajama v1, and peS2o  [149, 160]. We compare the aforementioned approaches in Table 4 and find that fastText-based filtering outperforms all other approaches. We next aim to understand how fastText training recipes affect its effectiveness as a data filtering network [55].
@@
To better understand the limits of fastText, we train several variants, exploring different choices for the reference data (i.e., the examples given positive labels), feature space, and filtering threshold, as shown in Table 5. For reference positive data, we considered commonly used sources like Wikipedia [59], OpenWebText2 [59], and RedPajama-books [160], following the reference data used for GPT-3 [30]. We also try a novel approach, using instruction-formatted data, drawing examples from OpenHermes 2.5 [157] (OH-2.5) and high-scoring posts from the r/ExplainLikeImFive (ELI5) subreddit. Overall, we find, when controlling for other hyperparameters, the fastText OH-2.5 +ELI5 approach gives a 3.5 percentage point lift on Core compared to the conventional choices. It is natural to ask whether using OH-2.5 data for filtering could preclude additional gains from instruction-tuning. In Appendix P, we show this is not the case, further suggesting the strength and compatibility of this approach with modern fine-tuning paradigms. Finally, we observe that using a fairly strict threshold, which keeps the top-10% of examples, helps over more permissive top-15% and top-20% thresholds. We further study the unintuitive behavior of dataset filtering and its connection to human judgment in Appendix M.
@@
Researchers often combine Common Crawl (CC) with other data sources that are considered high-quality [160, 59, 65, 162] (e.g., Wikipedia, arXiv, Stack exchange, and peS2o [149]). Since DCLM participants can include additional data sources in the bring your own data track, we examined the potential benefits of adding high-quality sources to training sets derived from Common Crawl only. We compare a model trained on 100% filtered CC data to models trained with the mixing proportion from Llama 1 and RedPajama: 67% CC, and 33% from Wikipedia, Books, Stack exchange, arXiv, and Github. For the CC component, we consider different variants: a subset of our DCLM-baseline, RedPajama’s CC portion, RefinedWeb, and C4. The results in Table 6 show that mixing improves performance for the lower-performing CC subsets (C4, RedPajama-CC, and RefinedWeb). In the case of DCLM-baseline however, mixing actually hurts performance on average, which suggests it can be counterproductive given performant filtering. For additional mixing results, see Appendix L.
@@
C4 23.7 25.9 ($+2.2$) 12.5 13.3 ($+0.8$)
RPJ CC only 24.0 25.7 ($+1.7$) 12.1 13.5 ($+1.4$)
RefinedWeb 25.1 26.5 ($+1.4$) 12.9 13.1 ($+0.2$)
DCLM-baseline  [^1] The DCLM-baseline subset used here relied on an earlier version of our classifier and our pool. 31.1 29.5 ($-1.6$) 16.0 15.3 ($-0.7$)
@@
The results of this analysis can be seen in Table 7. We see that this removal of contaminated samples does not lead to a decrease in performance for our model. As such, we can see that our performance gains in MMLU are not caused by increased presence of MMLU in our dataset.
@@
We also apply the above removal strategy on Dolma-V1.7 [150] and FineWeb-Edu [100], to measure what the contamination differences are between DCLM-baseline and those datasets. The results can be seen in Table 8. We see that our DCLM-baseline has roughly similar contamination stats as other high performing datasets according to this analysis. We provide further contamination analysis that extends to our entire benchmark suite in Appendix N.
@@
Here, we test if datasets that perform well on the DCLM benchmark also maintain their strength with an order of magnitude more compute. To ensure our trained model is broadly useful, including for math and coding tasks, we combine our 3.8T DCLM-baseline with the StarCoder [90] and ProofPile2 [14] data to arrive at a 4.1T token dataset. We train a 7B model for 2.5T tokens on this dataset with the same hyperparameters as our largest competition scale except for two separate cool-downs phase for the 200B and 270B tokens on a modified distribution that was 70% DCLM-baseline with a tighter fastText threshold, and 30% math datasets (see Appendix P). We then take a “model soup" of these two separate cool-downs[173]. We then adopt the continual pre-training methodology from Pouransari et al. [126] for 100B tokens on the same distribution to increase the context length from 2048 to 8192, we provide more details on this procedure in Section P.2.
In Table 9, we show that our model outperforms all 7B models trained on public training sets and approaches closed-data models trained for more tokens such as Llama-8B, Mistral-7B, and Gemma-7B. Additionally, in Appendix O, we show that our model achieves strong instruction-tuning performance. After instruction tuning on publicly available IT datasets, our model maintains most of its benchmark performance and achieves an AlpacaEval2.0 LC Win-rate of 16.6, which outperforms Gemma-Instruct (10.4), while approaching the strong performance of Mistral-v0.2-7B (17.1) and Llama3-Instruct (22.9).